# Diffusion & Adversarial Schrödinger Bridges via Iterative Proportional Markovian Fitting

### v4

@@ 4 Experimental Illustrations
Experimental setups. We consider multivariate Gaussian distributions for which we have closed-form IPMF update formulas, an illustrative 2D example, the Schrödinger Bridges Benchmark (gushchin2023building) and real-life image data distributions, i.e., the colored MNIST dataset and the CelebA dataset (liu2015faceattributes). All technical details can be found in the Appx. E.
@@
Remark. Notably, in practice, both the IPF and IMF procedures can be recovered through different implementations. For example, IPF can be realized through (D/A)SBM with the IPF starting coupling, or alternatively via DSB (de2021diffusion). IMF, in turn, can be implemented using (D/A)SBM with a one-directional parametrization. However, in practice, matching-based methods exhibit superior performance (shi2023diffusion). Furthermore, the authors of (shi2023diffusion; peluchetti2023diffusion; gushchin2024adversarial) observed that bidirectional IMF does not accumulate approximation errors, whereas relying solely on one direction parametrization leads to error accumulation and eventual divergence (de2024schr, Appx. I). Therefore, we argue that a direct comparison between the IPMF procedure and previous practical implementations is unnecessary.
@@ 4.3 Evaluation on the SB Benchmark
The best metric is bolded. Results marked with $\dagger$ are taken from (gushchin2024adversarial) and (gushchin2023building). The results of DSBM and ASBM algorithms starting from different starting processes are denoted as (D/A)SBM-*name of starting process*
@@ 4.4 Unpaired Image-to-Image translation
(a) $x\sim p_{0}$}
(b) DSBM-IMF-OT
(c) DSBM-Identity
(d) DSBM-SD SDEdit
(e) DSBM-DDPM SDEdit
(f) $x\sim p_{0}$}
(g) ASBM-IMF-OT
(h) ASBM-Identity
(i) ASBM-SD SDEdit
(j) ASBM-DDPM SDEdit
@@
[…]
DSBM ASBM
IMF SD SDEdit DDPM SDEdit Identity IMF DDPM SDEdit SD SDEdit Identity
@@ Appendix A Discussion
Potential impact. The IPMF procedure demonstrates a potential to overcome the error accumulation problem observed in distillation methods—such as rectified flows (liu2022flow; liu2023instaflow)—which are used to accelerate foundational image generation models like StableDiffusion 3 in (esser2024scaling). These distillation methods are the limit of one-directional IMF procedure with $\epsilon\rightarrow 0$. The one-directional version accumulates errors, which may lead to the divergence (de2024schr, Appx. I). The use of the bidirectional version (with $\epsilon>0$) should correct the marginals and make diffusion trajectories straighter to accelerate the inference of diffusion models. We believe that considering such distillation techniques from the IPMF perspective may help to overcome the current limitations of these techniques.
Another potential impact of our contribution is the advancement of multi-marginal SB methods. This direction has been explored only rarely in the literature (chen2019multi; chen2023deep; shen2025multi; howard2025schr; lavenant2021towards; theodoropoulos2025momentummultimarginalschrodingerbridge), mainly because the multi-marginal case is inherently difficult: it requires solving multiple two-marginal (classical) SB instances. A notable examples are (howard2025schr; theodoropoulos2025momentummultimarginalschrodingerbridge), which extends the IMF procedure to the multi-marginal setting. Within this context, our framework provides a way to select a suitable starting coupling for initialization, thereby offering a potential route to reducing the training burden. In this sense, our contribution may encourage more systematic and deeper analysis of multi-marginal SB.
@@ Appendix B Motivation for SB over Foundational Models
At first glance, one might consider foundational models as potential baselines, since translation via large text-to-image models trained on extensive image corpora may work adequately on the benchmark datasets we consider (CelebA, MNIST). However, they do not constitute a relevant baseline for the unpaired translation task, because their training data may lack the domain-specific examples required. In contrast, methods for solving the unpaired translation (including the SB) are designed to address domain‑specific tasks across various scientific fields (schneider2022whole; singh2024data; shi2023diffusion). Moreover, these methods successfully address non-image-related downstream tasks such as single-cell data analysis (tong2024simulation, Section 6), where large text-to-image models are just irrelevant.
@@ C.1 Schrödinger Bridge (SB) Problem in continuous-time
where $\mathcal{F}(p_{0},p_{1})\subset\mathcal{P}(C([0,1]),\mathbb{R}^{D})$ is the set of all stochastic processes pinned by marginal distributions $p_{0}$ and $p_{1}$ at times 0 and $1$, respectively. The minimization problem (13) has a unique solution $T^{*}$ which can be represented as forward or backward diffusion (leonard2013survey):
@@
It has been proved (leonard2013survey) that for the solution $T^{*}$ it’s conditional process is given by $T^{*}_{|x_{0},x_{1}}=W^{\epsilon}_{|x_{0},x_{1}}$. Thus, we can set $T_{|x_{0},x_{1}}=W^{\epsilon}_{|x_{0},x_{1}}$ zeroing the second term in (14) and minimize over processes with $T_{|x_{0},x_{1}}=W^{\epsilon}_{|x_{0},x_{1}}$. This leads to the equivalent Static formulation of the Schrödinger Bridge problem:
@@ C.2 Iterative Proportional Fitting (IPF) for continuous-time
As in the discrete-time case, these projections replace marginal distributions $p^{T}(x_{1})$ and $p^{T}(x_{0})$ in the processes $p^{T}(x_{1})T_{|x_{1}}$ and $p^{T}(x_{0})T_{|x_{0}}$ by $p_{1}(x_{1})$ and $p_{0}(x_{0})$ respectively. Similarly to discrete-time formulation, the sequence of $T^{k}$ converges to the solution of the Schrödinger Bridge problem $T^{*}$ implicitly decreasing the reverse Kullback-Leibler divergence $\text{KL}(T^{k}||T^{*})$ between the current process $T^{k}$ and the solution to the SB problem TT*. Additionally, it should be mentioned that existing methods perform projections via numerical approximation of forward and time-reversed conditional processes, $T_{|x_{0}}$ and $T_{|x_{1}}$, by learning their drifts via one of the methods: score matching (de2021diffusion) or maximum likelihood estimation (vargas2021solving).
@@ C.3 Iterative Markovian Fitting (IMF) for continuous-time
where we denote by $T_{M}$ the Markovian projections of the processes $T$, which can be represented as the forward or backward time diffusion as follows (gushchin2024adversarial, Section 2.1):
@@
This procedure converges to a unique solution, which is the Schrödinger bridge $T^{*}$ (leonard2013survey). While reciprocal projection can be easily done by combining the joint distribution $p^{T}(x_{0},x_{1})$ of the process $T$ and Brownian bridge $W^{\epsilon}_{|x_{0},x_{1}}$, the Markovian projection is much more challenging and must be fitted via Bridge matching (shi2023diffusion; liu2022let; peluchetti2023non).
@@ D.1 Gaussian plans as entropic optimal transport plans
Theorem 3.2 from gushchin2023building states that if the conditional distribution $q(x_{1}|x_{0})$ can be expressed as:
@@ D.3 Discrete IMF step analysis: multidimensional case for large $\epsilon$}
Following (gushchin2024adversarial), we have an explicit formula for the reciprocal step. For any $0\leq i,j\leq N+1$, we have marginal covariance $\Sigma_{t_{i},t_{i}}$ at time moment $t_{i}$ and joint covariance $\Sigma_{t_{i},t_{j}}$ between time moments $t_{i}$ and $t_{j}$:
@@
For now, consider $B$ such that its singular values are positive and distinct (note that the set of such matrices is dense in $\mathbb{R}^{D\times D}$). Then the SVD map $B\mapsto(U,\Lambda,V)$ such that $B=U\Lambda V^{*}$ is differentiable at $B$ (see magnus2019matrix, Section 3.8.8), thus so is the polar decomposition map $B\mapsto(Q,S)$ such that $B=KC$, where $K$ is orthogonal and $C$ is PSD matrices. As
@@
Furthermore, $0<h(x)\leq 1$ and $(xh(x^{2}))^{\prime}=\frac{2}{(1+\sqrt{1+4x^{2}})\sqrt{1+4x^{2}}}\in(0,1]$, hence $0\prec h(S^{2})\preccurlyeq I$ and $Ch(C^{2})$ is $1$-Lipschitz w.r.t. the Frobenius norm (wihler2009holder, Thm. 1.1). Note that $\|KdC\|_{F}^{2}=\operatorname{Tr}[(KdC)^{\top}dKC]=\operatorname{Tr}[(CdC)(K^{\top}dK)]=0$ since $K^{\top}dK$ is skew-symmetric. It can be shown from the orthogonality of $K$:
@@ D.5 IMF step analysis in $1D$
Continuous case. Following (peluchetti2023diffusion, Eq. 42), we have the formula for $\tilde{\rho}_{n}$:
@@
Discrete case ($N=1$). We use explicit formula (44) for a new correlation $\tilde{\rho}_{n}=f(\rho_{n})$ after D-IMF step from (gushchin2024adversarial) provided in the beginning of Section D.3.
@@
D.7 Proof of discrete IPMF General Convergence Theorem 3.3
@@
$\geq\alpha\int proj_{\mathcal{M}}[q](x_{1}|x_{t_{1}})d\mu(x_{t_{1}})=:\alpha\mu^{\prime}(x_{1}).$
@@
Finally, we argue by contradiction: if $q^{4k}_{0,1}\stackrel{{\scriptstyle w}}{{\nrightarrow}}q^{*}_{0,1}$, we can select a weakly converging subsequence $q^{4l_{i}}_{0,1}\xrightarrow[]{w}q^{\prime}_{0,1}\neq q^{*}_{0,1}$. But by (78) $q^{\prime}_{0,1}=q^{*}_{0,1}$. This finishes the proof. ∎

### v5

@@ 4 Experimental Illustrations
Experimental setups. We consider multivariate Gaussian distributions for which we have closed-form IPMF update formulas, an illustrative 2D example, the Schrödinger Bridges Benchmark (Gushchin et al., 2023b) and real-life image data distributions, i.e., the colored MNIST dataset and the CelebA dataset (Liu et al., 2015b). All technical details can be found in the Appx. E.
@@
Remark. Notably, in practice, both the IPF and IMF procedures can be recovered through different implementations. For example, IPF can be realized through (D/A)SBM with the IPF starting coupling, or alternatively via DSB (De Bortoli et al., 2021). IMF, in turn, can be implemented using (D/A)SBM with a one-directional parametrization. However, in practice, matching-based methods exhibit superior performance (Shi et al., 2023). Furthermore, the authors of (Shi et al., 2023; Peluchetti, 2023a; Gushchin et al., 2024b) observed that bidirectional IMF does not accumulate approximation errors, whereas relying solely on one direction parametrization leads to error accumulation and eventual divergence (De Bortoli et al., 2024, Appx. I). Therefore, we argue that a direct comparison between the IPMF procedure and previous practical implementations is unnecessary.
@@ 4.3 Evaluation on the SB Benchmark
The best metric is bolded. Results marked with $\dagger$ are taken from (Gushchin et al., 2024b) and (Gushchin et al., 2023b). The results of DSBM and ASBM algorithms starting from different starting processes are denoted as (D/A)SBM-*name of starting process*
@@ 4.4 Unpaired Image-to-Image translation
(a) $x_{0}\!\sim\!p_{0}\!\!\!$}
(b) DSBM-IMF-OT
(c) ASBM-IMF-OT
(d) $x_{1}\!\sim\!q^{0}\!$}
(e) $x_{0}\!\sim\!p_{0}\!\!$}
(f) DSBM-DDPM SDEdit
(g) ASBM-DDPM SDEdit
(h) $x_{1}\!\sim\!q^{0}\!$}
(i) $x_{0}\!\sim\!p_{0}\!\!\!$}
(j) DSBM-SD SDEdit
(k) ASBM-SD SDEdit
(l) $x_{1}\!\sim\!q^{0}$}
(m) $x_{0}\!\sim\!p_{0}\!\!\!$}
(n) DSBM-Identity
(o) ASBM-Identity
(p) $x_{1}\!\sim\!q^{0}\!$}
[…]
@@
Initialisation (coupling) DSBM ASBM
IMF DDPM SDEdit SD SDEdit Identity IMF DDPM SDEdit SD SDEdit Identity IMF DDPM SDEdit SD SDEdit Identity
@@
[…]
(a) MSE($x_{0}$, $\hat{x}_{1}$) for (D/A)SBM
(b) FID for DSBM
(c) FID for ASBM
[…]
(a) Input AFHQ cat images
(b) DSBM-IMF-OT translation results.
(c) DSBM-Identity translation results.
[…]
DSBM-IMF-OT 53.42 0.085 0.591
DSBM-Identity 65.19 0.054 0.731
[…]
AFHQ. For AFHQ (Choi et al., 2020), we consider classes cat and wild with $512\times 512$ resolution images. Each class contains approximately 5000 samples. We run DSBM with IMF-OT and Identity couplings and present the results in Figure 6 and Table E.10. We observe similar quality-similarity tradeoff as Celeba setup. We provide technical details for this setup in Appendix E.10.
@@ 5 Broader impact
Acknowledgements.
We thank Kirill Sokolov for the careful feedback and support in refining the proofs. The work was supported by the grant for research centers in the field of AI provided by the Ministry of Economic Development of the Russian Federation in accordance with the agreement  [^00] C313925P4F0002 and the agreement №139-10-2025-033.
@@ Appendix A Discussion
Potential impact. The IPMF procedure demonstrates a potential to overcome the error accumulation problem observed in distillation methods—such as rectified flows (Liu et al., 2022; 2023b)—which are used to accelerate foundational image generation models like StableDiffusion 3 in (Esser et al., 2024). These distillation methods are the limit of one-directional IMF procedure with $\epsilon\rightarrow 0$. The one-directional version accumulates errors, which may lead to the divergence (De Bortoli et al., 2024, Appx. I). The use of the bidirectional version (with $\epsilon>0$) should correct the marginals and make diffusion trajectories straighter to accelerate the inference of diffusion models. We believe that considering such distillation techniques from the IPMF perspective may help to overcome the current limitations of these techniques.
Another potential impact of our contribution is the advancement of multi-marginal SB methods. This direction has been explored only rarely in the literature (Chen et al., 2019; 2023a; Shen et al., 2025; Howard et al., 2025; Lavenant et al., 2021; Theodoropoulos et al., 2025), mainly because the multi-marginal case is inherently difficult: it requires solving multiple two-marginal (classical) SB instances. A notable examples are (Howard et al., 2025; Theodoropoulos et al., 2025), which extends the IMF procedure to the multi-marginal setting. Within this context, our framework provides a way to select a suitable starting coupling for initialization, thereby offering a potential route to reducing the training burden. In this sense, our contribution may encourage more systematic and deeper analysis of multi-marginal SB.
@@
Method View of bidirectional procedure Convergence guarantees Starting coupling
DSBM (Shi et al., 2023) A heuristic approach for mitigating error accumulation OnlyforonedirectionalcontinuoustimeprocedurewithIMFandIPFstartingcouplingsOnlyforone-directionalcontinuous-timeprocedurewithIMFandIPFstartingcouplings Only IMF and IPF
ASBM (Gushchin et al., 2024b) A heuristic approach for mitigating error accumulation OnlyforonedirectionaldiscretetimeprocedurewithIMFstartingcouplingsOnlyforone-directionaldiscrete-timeprocedurewithIMFstartingcouplings Only IMF
[…]
Table 3: Positioning of our IPMF framework relative to prior bidirectional SB heuristics.
@@ Appendix B Motivation for SB over Foundational Models
At first glance, one might consider foundational models as potential baselines, since translation via large text-to-image models trained on extensive image corpora may work adequately on the benchmark datasets we consider (CelebA, MNIST). However, they do not constitute a relevant baseline for the unpaired translation task, because their training data may lack the domain-specific examples required. In contrast, methods for solving the unpaired translation (including the SB) are designed to address domain‑specific tasks across various scientific fields (Schneider et al., 2022; Singh et al., 2024; Shi et al., 2023). Moreover, these methods successfully address non-image-related downstream tasks such as single-cell data analysis (Tong et al., 2024, Section 6), where large text-to-image models are just irrelevant.
@@ C.1 Schrödinger Bridge (SB) Problem in continuous-time
where $\mathcal{F}(p_{0},p_{1})\subset\mathcal{P}(C([0,1]),\mathbb{R}^{D})$ is the set of all stochastic processes pinned by marginal distributions $p_{0}$ and $p_{1}$ at times 0 and $1$, respectively. The minimization problem (13) has a unique solution $T^{*}$ which can be represented as forward or backward diffusion (Léonard, 2013):
@@
It has been proved (Léonard, 2013) that for the solution $T^{*}$ it’s conditional process is given by $T^{*}_{|x_{0},x_{1}}=W^{\epsilon}_{|x_{0},x_{1}}$. Thus, we can set $T_{|x_{0},x_{1}}=W^{\epsilon}_{|x_{0},x_{1}}$ zeroing the second term in (14) and minimize over processes with $T_{|x_{0},x_{1}}=W^{\epsilon}_{|x_{0},x_{1}}$. This leads to the equivalent Static formulation of the Schrödinger Bridge problem:
@@ C.2 Iterative Proportional Fitting (IPF) for continuous-time
As in the discrete-time case, these projections replace marginal distributions $p^{T}(x_{1})$ and $p^{T}(x_{0})$ in the processes $p^{T}(x_{1})T_{|x_{1}}$ and $p^{T}(x_{0})T_{|x_{0}}$ by $p_{1}(x_{1})$ and $p_{0}(x_{0})$ respectively. Similarly to discrete-time formulation, the sequence of $T^{k}$ converges to the solution of the Schrödinger Bridge problem $T^{*}$ implicitly decreasing the reverse Kullback-Leibler divergence $\text{KL}(T^{k}||T^{*})$ between the current process $T^{k}$ and the solution to the SB problem TT*. Additionally, it should be mentioned that existing methods perform projections via numerical approximation of forward and time-reversed conditional processes, $T_{|x_{0}}$ and $T_{|x_{1}}$, by learning their drifts via one of the methods: score matching (De Bortoli et al., 2021) or maximum likelihood estimation (Vargas et al., 2021).
@@ C.3 Iterative Markovian Fitting (IMF) for continuous-time
where we denote by $T_{M}$ the Markovian projections of the processes $T$, which can be represented as the forward or backward time diffusion as follows (Gushchin et al., 2024b, Section 2.1):
@@
This procedure converges to a unique solution, which is the Schrödinger bridge $T^{*}$ (Léonard, 2013). While reciprocal projection can be easily done by combining the joint distribution $p^{T}(x_{0},x_{1})$ of the process $T$ and Brownian bridge $W^{\epsilon}_{|x_{0},x_{1}}$, the Markovian projection is much more challenging and must be fitted via Bridge matching (Shi et al., 2023; Liu et al., ; Peluchetti, 2023b).
@@ D.1 Gaussian plans as entropic optimal transport plans
Theorem 3.2 from Gushchin et al. (2023b) states that if the conditional distribution $q(x_{1}|x_{0})$ can be expressed as:
@@ D.3 Discrete IMF step analysis: multidimensional case for large $\epsilon$}
Following (Gushchin et al., 2024b), we have an explicit formula for the reciprocal step. For any $0\leq i,j\leq N+1$, we have marginal covariance $\Sigma_{t_{i},t_{i}}$ at time moment $t_{i}$ and joint covariance $\Sigma_{t_{i},t_{j}}$ between time moments $t_{i}$ and $t_{j}$:
@@
For now, consider $B$ such that its singular values are positive and distinct (note that the set of such matrices is dense in $\mathbb{R}^{D\times D}$). Then the SVD map $B\mapsto(U,\Lambda,V)$ such that $B=U\Lambda V^{*}$ is differentiable at $B$ (see Magnus and Neudecker, 2019, Section 3.8.8), thus so is the polar decomposition map $B\mapsto(Q,S)$ such that $B=KC$, where $K$ is orthogonal and $C$ is PSD matrices. As
@@
Furthermore, $0<h(x)\leq 1$ and $(xh(x^{2}))^{\prime}=\frac{2}{(1+\sqrt{1+4x^{2}})\sqrt{1+4x^{2}}}\in(0,1]$, hence $0\prec h(S^{2})\preccurlyeq I$ and $Ch(C^{2})$ is $1$-Lipschitz w.r.t. the Frobenius norm (Wihler, 2009, Thm. 1.1). Note that $\|KdC\|_{F}^{2}=\operatorname{Tr}[(KdC)^{\top}dKC]=\operatorname{Tr}[(CdC)(K^{\top}dK)]=0$ since $K^{\top}dK$ is skew-symmetric. It can be shown from the orthogonality of $K$:
@@ D.5 IMF step analysis in $1D$
Continuous case. Following (Peluchetti, 2023a, Eq. 42), we have the formula for $\tilde{\rho}_{n}$:
@@
Discrete case ($N=1$). We use explicit formula (44) for a new correlation $\tilde{\rho}_{n}=f(\rho_{n})$ after D-IMF step from (Gushchin et al., 2024b) provided in the beginning of Section D.3.
@@
D.7 Proof of IPMF General Convergence Theorem 3.3
@@
We split the proof into two parts. First, consider the discrete case.
Discrete case.
@@
$\geq\alpha\int proj_{\mathcal{M}}[q](x_{1}|x_{t_{1}})d\mu(x_{t_{1}})=:\alpha\mu^{\prime}(x_{1}).$ (76)
@@
Finally, we argue by contradiction: if $q^{4k}_{0,1}\stackrel{{\scriptstyle w}}{{\nrightarrow}}q^{*}_{0,1}$, we can select a weakly converging subsequence $q^{4l_{i}}_{0,1}\xrightarrow[]{w}q^{\prime}_{0,1}\neq q^{*}_{0,1}$. But by (79) $q^{\prime}_{0,1}=q^{*}_{0,1}$. This finishes the proof.
Continuous case.
We now explain how to extend the above argument to the continuous-time setting. The key point is to verify a Doeblin minorization condition for the Markovian process obtained after the projection step (see, e.g., Section 2 in Stroock (2005)).
Fix some $\delta\in(0,1/2)$. For each $k\in\mathbb{N}$, let $(X_{t}^{4k+2})_{t\in[0,1]}$ denote the Markov diffusion corresponding to the law $q^{4k+2}$, and let
$P_{k}(x,A):=\mathbb{P}\bigl(X^{4k+2}_{1}\in A\,\big|\,X^{4k+2}_{0}=x\bigr),\qquad x\in\mathbb{R}^{D},\ A\in\mathcal{B}(\mathbb{R}^{D}),$
be its transition kernel from time 0 to time $1$. This admits a decomposition of the evolution on $[0,1]$ into three subintervals $0,\delta]$, $\delta,1-\delta]$ and $1-\delta,1]$.
Disspaptivity. Next, we show that, by construction of the Markovian projection, the drift on $\delta,1-\delta]$ is Lipschitz and dissipative. Set
$b^{+}(t,x):=\mathbb{E}\!\left[\frac{X_{1}-x}{1-t}\,\middle|\,X_{t}=x\right]=\frac{m_{t}(x)-x}{1-t},\qquad m_{t}(x):=\mathbb{E}[X_{1}\mid X_{t}=x],$
and recall that since the supports of $p_{0}(x)$ and $p_{1}(x)$ are bounded, $\lVert X_{1}\rVert\leq M_{1}$ and $\|m_{t}(x)\|\leq M_{1}$ for any $x$. Then
$\langle b^{+}(t,x),x\rangle=\frac{\langle m_{t}(x),x\rangle-\|x\|^{2}}{1-t}\leq\frac{M_{1}\|x\|-\|x\|^{2}}{1-t}\leq-\frac{1}{2(1-t)}\|x\|^{2}+\frac{M_{1}^{2}}{2(1-t)}.$
This yields uniform coercivity (Lyapunov-type dissipativity) for any $t\in[\delta,1-\delta]$.
The drift is Lipshitz. Consider
$\nabla_{x}b^{+}(t,x)=\frac{\nabla_{x}m_{t}(x)-I}{1-t},$
and recall that the transition density is
$\pi_{x}(x_{1}):=q_{t,1}(x_{1}\mid x)\propto g(x_{1})\exp\left(-\frac{\|x_{1}-x\|^{2}}{2\varepsilon^{2}(1-t)}\right).$
Consequently,
$\nabla_{x}m_{t}(x)$ $=\nabla_{x}\mathbb{E}[X_{1}|X_{t}=x]=\nabla_{x}\int x_{1}\pi_{x}(x_{1})d\,x_{1}=\int x_{1}\nabla_{x}\pi_{x}(x_{1})\,dx_{1}$
$=\int(x_{1}\times\nabla_{x}\log\pi_{x}(x_{1}))\pi_{x}(x_{1})\,dx_{1}=\frac{1}{2\varepsilon^{2}(1-t)}\mathrm{Cov}_{\pi_{x}}[X_{1}].$
Thus,
$\lVert\nabla_{x}m_{t}(x)\rVert\leq C\frac{\mathrm{diam}^{2}(\mathrm{supp}\,p_{1})}{\varepsilon^{2}(1-t)}.$
This ensures
$\|\nabla_{x}b^{+}(t,x)\|\leq\frac{1}{1-t}+C\frac{\mathrm{diam}(\mathrm{supp}\,p_{1})^{2}}{\varepsilon^{2}(1-t)^{2}}.$
A similar result holds for the backward drift
$b^{-}(t,x):=\mathbb{E}\!\left[\frac{x-X_{0}}{t}\,\middle|\,X_{t}=x\right]=\frac{x-m^{\prime}_{t}(x)}{t},\qquad m^{\prime}_{t}(x):=\mathbb{E}[X_{0}\mid X_{t}=x].$
Small-set condition on $\delta,1-\delta]$. Fix $R>0$ and consider the transition from time $\delta$ to $1-\delta$. Denote by $p_{k}(x,y)$ the transition density of $X^{4k+2}_{1-\delta}$ given $X^{4k+2}_{\delta}=x$. Since the diffusion coefficient is constant non-degenerate ($\varepsilon^{2}I_{D}$) and the drift is locally Lipschitz (with a uniform bound on $\|\nabla b^{+}(\cdot)\|$ on $\delta,1-\delta]$), heat-kernel lower bounds for uniformly elliptic diffusions yield that $p_{k}$ is continuous and strictly positive, and admits a lower bound on compact sets with constants independent of $k$; see, e.g., Theorem 7 in Aronson (1968).
In particular, for any fixed $r>0$ there exists $m_{R,r}>0$ such that
$\inf_{k\in\mathbb{N}}\ \inf_{x\in B_{R},\ y\in B_{r}}p_{k}(x,y)\ \geq\ m_{R,r}\ >0.$
Let $\nu_{r}$ be the uniform probability measure on $B_{r}$ and set $\beta_{R}:=m_{R,r}\,\mathrm{Leb}(B_{r})$. Then for all $k\in\mathbb{N}$, all $x\in B_{R}$, and all measurable $A$,
$\mathbb{P}\bigl(X^{4k+2}_{1-\delta}\in A\,\big|\,X^{4k+2}_{\delta}=x\bigr)\ \geq\ \beta_{R}\,\nu_{r}(A)$ (80)
By the Markov property, for any measurable $A$,
$\mathbb{P}\bigl(X^{4k+2}_{1}\in A\,\big|\,X^{4k+2}_{\delta}=x\bigr)=\int\mathbb{P}\bigl(X^{4k+2}_{1}\in A\,\big|\,X^{4k+2}_{1-\delta}=z\bigr)\,\mathbb{P}\bigl(X^{4k+2}_{1-\delta}\in dz\,\big|\,X^{4k+2}_{\delta}=x\bigr).$
Combining this with (80), we obtain
$\mathbb{P}\bigl(X^{4k+2}_{1}\in A\,\big|\,X^{4k+2}_{\delta}=x\bigr)\geq\beta_{R}\mu_{r,k}(A),\qquad x\in B_{R},$
where $\mu_{r,k}$ is the probability measure defined by
$\color[rgb]{0,0,0}\mu_{r,k}(A):=\int\mathbb{P}\bigl(X^{4k+2}_{1}\in A\,\big|\,X^{4k+2}_{1-\delta}=z\bigr)\,\nu_{r}(dz).$} (81)
Next, we control the tails of distribution of $X^{4k+2}_{\delta}$ uniformly in $k$. By the definition of the reciprocal projection, the segment $[0,1]$ between $X_{0}$ and $X_{1}$ is (conditionally on $(X_{0},X_{1})$) distributed as a Brownian bridge with variance parameter $\sigma^{2}=\varepsilon^{2}\delta(1-\delta)$. Hence, the marginal at time $\delta$ is a mixture of Gaussian laws with covariance matrix $\sigma^{2}I_{D}$ and mean
$m_{\delta}(x_{0},x_{1})=(1-\delta)x_{0}+\delta x_{1},$
where $(x_{0},x_{1})$ ranges over the support of the endpoint coupling. Since the supports of $p_{0}$ and $p_{1}$ are bounded, there exists $R_{0}>0$ such that $\|m_{\delta}(x_{0},x_{1})\|\leq R_{0}$ for all $(x_{0},x_{1})$ in this support. Standard Gaussian tail bounds then imply that, for any $\eta\in(0,1)$, we can choose $R>0$ large enough so that
$\sup_{k\in\mathbb{N}}\ \sup_{x\in\mathrm{supp}(p_{0})}\mathbb{P}\bigl(X^{4k+2}_{\delta}\notin B_{R}\,\big|\,X^{4k+2}_{0}=x\bigr)\;\leq\;\eta.$
Equivalently,
$\mathbb{P}\bigl(X^{4k+2}_{\delta}\in B_{R}\,\big|\,X^{4k+2}_{0}=x\bigr)\;\geq\;1-\eta,\qquad x\in\mathrm{supp}(p_{0}),\ k\in\mathbb{N}.$ (82)
Combining (81) and (82), we obtain, for $x\in\mathrm{supp}(p_{0})$ and any measurable $A\subset\mathbb{R}^{D}$,
$P_{k}(x,A)$ $=\mathbb{E}\Bigl[\mathbb{P}\bigl(X^{4k+2}_{1}\in A\,\big|\,X^{4k+2}_{\delta}\bigr)\,\Big|\,X^{4k+2}_{0}=x\Bigr]$
$\geq\mathbb{E}\Bigl[\mathbb{P}\bigl(X^{4k+2}_{1}\in A\,\big|\,X^{4k+2}_{\delta}\bigr)\mathbf{1}_{\{X^{4k+2}_{\delta}\in B_{R}\}}\,\Big|\,X^{4k+2}_{0}=x\Bigr]$
$\geq\beta_{R}\nu_{R}(A)\,\mathbb{P}\bigl(X^{4k+2}_{\delta}\in B_{R}\,\big|\,X^{4k+2}_{0}=x\bigr)$
$\geq\beta_{R}(1-\eta)\,\mu_{r,k}(A).$
Thus, for all $x\in\mathrm{supp}(p_{0})$ and all $k\in\mathbb{N}$,
$P_{k}(x,\cdot)\;\geq\;\alpha\,\mu_{r,k}(\cdot)\quad\text{with}\quad\alpha:=\beta_{R}(1-\eta)\in(0,1),\ \ \mu:=\nu_{R}.$
That is, the family of kernels $(P_{k})_{k}$ satisfies a uniform Doeblin minorization on $\mathrm{supp}(p_{0})$.
It is well known that such a minorization implies total-variation contraction: for any probability measures $\lambda,\lambda^{\prime}$ on $\mathrm{supp}(p_{0})$,
$\|\lambda P_{k}-\lambda^{\prime}P_{k}\|_{\mathrm{TV}}\;\leq\;(1-\alpha)\,\|\lambda-\lambda^{\prime}\|_{\mathrm{TV}},\qquad k\in\mathbb{N}.$
Applying this with $\lambda=q^{4k+1}_{0}$ and $\lambda^{\prime}=p_{0}$ yields
$\|q^{4k+1}_{1}-p_{1}\|_{\mathrm{TV}}\;=\;\|q^{4k+1}_{0}P_{k}-p_{0}P_{k}\|_{\mathrm{TV}}\;\leq\;(1-\alpha)\,\|q^{4k+1}_{0}-p_{0}\|_{\mathrm{TV}}.$
The convergence $q^{4k+1}_{0}\to_{\mathrm{TV}}p_{0}$ is shown by the same argument applied backward in time (interchanging the roles of $p_{0}$ and $p_{1}$), and we conclude that
$q^{4k+1}_{1}\xrightarrow[k\to\infty]{\mathrm{TV}}p_{1}.$
In particular, the continuous-time analogue of (77) holds.
Having established the convergence of the marginals via the Doeblin minorization, it remains to ensure the well-posedness and stability of the iterative procedure. Next, we recall that the drift $b^{\pm}(t)$ is locally Lipschitz and dissipative.
Next, by the symmetry of the IMF procedure, it suffices to consider the transitions to the midpoint, $(0,1/2)$ and $(1,1/2)$. By standard diffusion theory, the regularity of the drift ensures that the corresponding Markovian transition kernels are smooth. Consequently, because the updated joint distribution $q^{i+1}(x_{0},x_{1/2},x_{1})$ is constructed by integrating these transition kernels against the previous distribution $q^{i}$, the mapping $q^{i}\mapsto q^{i+1}$ preserves weak convergence. Therefore, the IMF iteration is continuous with respect to the weak topology. The rest of the proof is similar to the discrete case. ∎